# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v2

@@ VI Results
TABLE I: Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (crash or accuracy), and the models in which the error was observed, along with their associated opsets (op.).
@@
1 lift_lexical_references Removes used value references. Crash EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init List index out of range. Crash YOLOv3 (op. 10, 12)
3 split_init split_predict Field is required but missing. Crash DenseNet121 (op. 7-9), YOLOv2 (op. 10), SSD (op. 12), Tiny-YOLOv3 (op. 11) YOLOv3 (op. 10, 12)
4 split_predict Required inputs are missing. Crash SSD (op. 12), DenseNet121 (op. 7), InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash YOLOv3 (op. 10), DenseNet121 (op. 7) InceptionV2 (op. 6-9)
6 split_predict Input element number mismatch. Crash FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather Indices value not equal to node dims size Crash Faster R-CNN: R-50-FPN (op. 10)
8 All Passes Implicit ONNX format version change. Accuracy Models of cases $1$-$11$ in Table II
9 rewrite_input_dtype Output name used multiple times. Crash GPT-2 (op. 10)
10 eliminate_ consecutive_ idempotent_ops Incompatible shape error. Crash (Run) Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12), R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11 eliminate_ duplicate_ initializer Erase output assertion failed. Crash Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv Duplicate initializer introduced. Crash (Run) RetinaNet (op. 9)
13 eliminate_shape_gather eliminate_nop_reshape Added initializers not used by the model. Warning Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash VGG-16-BN (op. 7), VGG-19-BN (op. 7), ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Accuracy Efficient-Lite4 (op. 11), SSD (op. 12)
VI-A Optimizer Crashes
Out of the $15$ detected bugs, $12$ bugs were related to model crashes, $2$ bugs resulted in output differences between the original and the optimized model, and $1$ bug introduced redundant initializers in the model that produced runtime warnings. In particular, we observed crashes were caused by $8$ out of the $47$ passes ($17$%), and occurred for specific models, with object detection models being more prone to errors from these $8$ passes. In particular, R-CNN models, such as Faster R-CNN R-50-FPN [54], RetinaNet [36], VGG-16, VGG-19 [58], as well as YOLOv3 [53], presented crashes by using the optimizer, in a range of opset instances and in passes related to the modification of gather nodes, and the elimination of idempotent operations. In total, we observed $31$ model instances where either the optimizer crashed upon processing them, or the optimized model was problematic, leading to a runtime crash. Some secondary passes are known to be unstable (e.g., split_init and split_predict describe in their documentation that they not strictly comply with the ONNX format), and therefore we do not focus on them as sources of faults in the optimization process. Overall, $12$ out of the $31$ model instances ($9.2$%) that were associated with a crash occurred while using passes related to fuse and eliminate operations, which are part of the default bundle of optimizations in the ONNX Optimizer and are therefore of particular significance.
As shown in Table I, crash faults can be attributed to three main causes: (1) model input issues (cases #$4$ and #$6$), (2) incorrect use of initializers (cases #$12$ and #$13$), and (3) issues associated with node properties such as duplicate names, obsolete fields, and incorrect dimensions (all other cases). Overall, object detection models were more prone to optimizer crashes compared to classification models. Due to their increased structural complexity and larger size, they are more likely to exhibit errors related to node properties and initializers, which account for the vast majority of crash faults. In contrast, only one issue was observed in the most complex model category — text generation models, that was related to model input in GPT-2. However, upon further inspection, we found that the transformation passes of the ONNX Optimizer had limited impact on text generation models — the majority of optimization passes left the models unchanged, and those that did make modifications affected a small proportion of nodes with negligible effect on outputs.
@@ VI-A Optimizer Crashes
VI-B Output Label Differences in Optimized Classification & Object Detection Models
@@
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ ($30.1$%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ ($16.6$%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($<0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences primarily affected the top-5 and top-10 predictions, with no changes in the top-1 predictions, except for the SSD model, which showed a $2.9\%$ difference in top-1 prediction. Although top-1 accuracy is often prioritized, it is important to recognize that changes in lower-ranked predictions can also indicate issues. Our methodology identified these inconsistencies, which have been reported to the ONNX Optimizer developers.
Figure 2: Illustrative example of problematic label classification and object detection between original unoptimized (blue) and optimized (orange) models. The bounding box location is affected for both bird instances, while the object label prediction is affected by model optimization on just the right bird instance, changing the label from bird to jet.
@@
TABLE II: Output label accuracy differences across classification (CL) and object detection (OD) models, showing the type, opset versions, and accuracy differences for top-1, top-5, and top-10 output predictions, between the original and optimized (using the default ONNX Optimizer setting) versions of each model, along with the issue related to the observed differences.
# Model Type Opset(s) Top 1 Top 5 Top 10 Issue
1 AlexNet [30] CL 7-9 0% 0% 0.02% Format Version Change
2 DenseNet121 [25] CL 7-9 0% 0% 0.02% Format Version Change
3 GoogleNet [62] CL 7-9 0% 0% 0.04% Format Version Change
4 Inception-1 [61] CL 7-9 0% 0% 0.25% Format Version Change
5 Inception-2 [61] CL 7-9 0% 0.02% 0.05% Format Version Change
6 MobileNet v2 [24, 56] CL 7-9 0% 0% 0.02% Format Version Change
7 ResNet18 [22] CL 7 0% 0.02% 0.02% Format Version Change
8 ResNet34-v2 [23] CL 7 0% 0.02% 0.02% Format Version Change
9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
10 ShuffleNet-v1 [74] CL 7-9 0% 0.7% 1% Format Version Change
11 SqueezeNet 1.0 [26] CL 7-9 0% 0% 0.02% Format Version Change
12 EfficientNet-Lite4 [63] CL 11 0% 0.04% 1% Graph Malformation
13 SSD [39] OD 12 2.9% 4.3% 6% Graph Malformation
@@ 9 ResNet50-caffe2 [22] CL 7-9 0% 2% 13% Format Version Change
Implicit ONNX format version upgrade:
@@
The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication. This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1). Despite being outdated, these opset versions are still prevalent, representing approximately 41% of the models in the ONNX Model Hub. While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions. We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
Problematic fusion pass:
@@
The pass fuse_bn_to_conv fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNormalization nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node. We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model. In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula. This behavior is illustrated on the right side of Figure 3. This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions. Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided. We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
@@
Table II shows label differences in top-K outputs for classification and object detection models, highlighting the causes. As an example, Figure 2 illustrates how an optimized model (orange) can misclassify and alter bounding boxes compared to the original model (blue).
Figure 3: Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4. The nodes affected are highlighted. The result should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
@@
TABLE III: Object detection metric differences across the original and optimized versions of object detection models, using the default ONNX Optimizer setting. The table presets the name, opset, average F1 score, precision, recall, mean Average Precision (MAP) and average Intersection over Union (IoU scores), for different IoU thresholds ($50$%, $75$%, and $90$%).
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.28% 95.28% 95.28% 86.62% 97.37%
Tiny YOLOv3 [53] 11 99.56% 99.56% 99.56% 90.57% 99.61%
YOLOv2 [52] 9 99.95% 99.95% 99.95% 90.85% 88.50%
YOLOv3 [53] 10 98.83% 98.83% 98.83% 86.66% 98.78%
YOLOv3 [53] 12 98.83% 98.83% 98.83% 86.65% 98.78%
(a) IoU threshold of 50%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [53] 11 99.36% 99.36% 99.36% 89.97% 99.61%
YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [53] 10 97.87% 97.87% 97.87% 85.71% 98.78%
YOLOv3 [53] 12 97.87% 97.87% 97.87% 85.71% 98.78%
(b) IoU threshold of 75%.
Model Op. Avg. F1 Avg. Prec. Avg. Recall mAP Avg. IoU
SSD [39] 12 95.27% 95.27% 95.27% 86.6% 97.37%
Tiny YOLOv3 [53] 11 99.34% 99.34% 99.34% 89.97% 99.61%
YOLOv2 [52] 9 99.97% 99.97% 99.97% 90.87% 88.60%
YOLOv3 [53] 10 97.8% 97.8% 97.8% 85.69% 98.78%
YOLOv3 [53] 12 97.8% 97.8% 97.8% 85.69% 98.78%
(c) IoU threshold of 90%.
VI-C Bounding Box Differences in Object Detection Models
We observed notable differences in bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table III with the following main findings:
YOLOv3 Models:
Both YOLOv3 models (op.set $10$ and $12$) achieve high average IoU, F1 score, precision, and recall—all above $98$%. Although the optimized model occasionally misplaces bounding boxes, these errors do not cause misclassifications, as none were found in the top-10 label predictions.
YOLOv2 Model:
YOLOv2 shows lower IoU values, likely because it uses multiple bounding boxes per detected object. Small differences between the original and optimized models accumulate, lowering IoU. However, high F1, precision, and recall indicate that detection performance remains largely unaffected, as demonstrated by the bounding boxes corresponding to the bird on the left in Figure 2.
SSD Model:
In contrast, SSD exhibits a $4.8$% drop in F1, precision, and recall between original and optimized models, resulting in 2.92.92.92.96666% misclassification rates despite maintaining a high IoU ($>97$%). This behavior is illustrated in Figure 2 by the bounding boxes associated with the bird on the right.
All models show a substantial mAP difference of up to $14.31$% between original and optimized model versions. This suggests the optimizer may cause certain classes to be missed, affecting both object detection and semantic segmentation. Consequently, even if detected objects are classified correctly, some objects in the scene might be overlooked, which could significantly impact real-world applications.
VI-D Optimizer Warnings
@@
We observed two issues related to the addition of redundant elements, specifically unused initializer entries introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects. We have reported it to the optimizer developers for additional review.
VII Tool Usability
@@
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub [7], which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories [6], it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets [3] package or from local storage.
@@
VIII Threats To Validity
We list the external and internal threats to validity in our evaluation below.
VIII-A External Threats
There are three primary external threats to the validity of our findings.
(1) Model Selection: Our evaluation included 130 models sourced from the official ONNX Model Hub, covering widely used categories such as classification, object detection, and text comprehension and generation. These models were selected for their maturity, accessibility, and broad adoption across academic and industrial domains. While we aimed to create a representative and diverse experimental set, the selection is not exhaustive. Models with opset versions below 7 were excluded due to incompatibility with the current ONNX and ONNX Runtime versions, which may limit our findings to more recent models. Additionally, to keep the evaluation tractable, we did not include certain model types—such as diffusion models and reinforcement learning—which may behave differently under optimization. Although our tool supports any ONNX-compliant model hub, the exclusion of these model types presents a potential threat to the generalizability of our results. We plan to address this limitation in future work.
To keep the experiments tractable, we did not evaluate model types such as diffusion models and reinforcement learning in this work; however, we plan to include them in future studies.
(2) Dataset Selection: Our experimental evaluation relies on widely used, real-world benchmark datasets tailored to the specific tasks of the models under test. While this approach ensures relevance and practical applicability—e.g., using ILSVRC [55] for image classification and object detection models like ResNet [22] and GoogleNet [62], and leveraging established HuggingFace datasets [3] for NLP tasks—it may limit generalizability. The use of popular datasets could introduce bias toward well-studied scenarios and may not reflect performance on less common or emerging tasks.
(3) Devices Used: Our experiments were conducted on two hardware platforms: an Intel i5 and an Intel i7 processor. While we verified consistency by (a) randomly sampling models of each type and running them on both devices with consistent results, and (b) referencing prior work by Louloudakis et al. [41], which found that optimization does not affect accuracy across processing devices, this still presents a potential threat to validity. The limited range of hardware may not capture performance or accuracy variations that could arise on other architectures, such as GPUs or edge devices, and may limit the generalizability of our findings.
VIII-B Internal Threats
We identify the following internal threat to validity in our findings.
Optimization Pass Ordering: The order in which optimization passes are applied can significantly affect model performance and behavior, as highlighted in prior work [31]. This is a well-known challenge in compiler research. While this work does not explore the impact of pass ordering in depth, we adopted two evaluation modes in OODTE: (1) the default ONNX Optimizer configuration, which includes commonly used passes such as fuse and eliminate, and (2) a configuration where all 47 available optimization passes were applied atomically. Although this strategy does not account for potential issues arising from specific pass combinations or orderings, it has proven effective in uncovering correctness issues in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
IX Discussion

### v3

@@ 6.1. Optimizer Crashes
Table 2. Overview of the issues identified using OODTE, detailing the responsible optimization pass, the detected error, its type (Crash, Accuracy (Acc.), or Warning (Warn.)), and the models in which the error was observed, along with their opsets (op.).
@@
1 lift_lexical_references
Removes used
value references.
Crash
EfficientNet-Lite4 (op. 11), SSD (op. 12)
2 split_init
List index
out of range.
Crash YOLOv3 (op. 10, 12)
3
split_init
split_predict
Required Field
is missing.
Crash
DenseNet121 (op. 7-9), YOLOv2 (op. 10),
SSD (op. 12), Tiny-YOLOv3 (op. 11),
YOLOv3 (op. 10, 12)
4 split_predict
Required inputs
are missing.
Crash
SSD (op. 12), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
5 split_predict Bad spec for node. Crash
YOLOv3 (op. 10), DenseNet121 (op. 7),
InceptionV2 (op. 6-9)
6 split_predict
Mismatch in number
of input elements
Crash
FCN ResNet-50-int8 (op. 12)
7 eliminate_shape_gather
Indices value
not equal to
node dims size
Crash
Faster R-CNN: R-50-FPN (op. 10)
8 All Passes
Implicit ONNX
format version
update.
Acc.
Models of cases 11-$11$
in Table 3
9 rewrite_input_dtype
Output name
used multiple times.
Crash GPT-2 (op. 10)
10
eliminate_consecutive_idempotent_ops
Incompatible shape
error.
Crash
Mask R-CNN/Faster R-CNN: R-50-FPN-int8 (12),
R-50-FPN-qdq (op. 12), R-50-FPN-fp32 (op. 12)
11
eliminate_duplicate_initializer
Erase output
assertion failed.
Crash
Faster R-CNN: R-50-FPN-qdq (12)
12 fuse_bn_into_conv
Duplicate initializer
introduced.
Crash
RetinaNet (op. 9)
13
eliminate_shape_gather
eliminate_nop_reshape
Added initializers
not used by the model.
Warn.
Faster R-CNN: R-50-FPN-int8 (op. 12)
14 All Passes Invalid input scale. Crash
VGG-16-BN (op. 7), VGG-19-BN (op. 7),
ResNet101_DUC_HDC (op. 7)
15 fuse_bn_into_conv Bad graph structure. Acc.
Efficient-Lite4 (op. 11),
SSD (op. 12)
@@ 8 All Passes
6.2. Output Label Differences in Optimized Classification & Object Detection Models
@@
Out of the $130$ optimized models tested, we observed unexpected output differences between the original and optimized models. These differences manifested as (1) label differences for classification models (e.g., ResNet), (2) label and/or bounding box differences for object detection models (e.g., YOLOv3), and (3) textual output differences in token sequences or predicted spans for text generation tasks. Since the ONNX Optimizer transformations are graph-based and do not alter numeric precision, output differences are not expected. Among the $93$ classification models tested, $28$ (30.130.1%) exhibited discrepancies in output labels. Similarly, $5$ out of $30$ (16.616.6%) object detection models presented output differences. The text generation models, on the other hand, were robust with only one instance (GPT-2) exhibiting negligible differences ($0.001\%$) while all others showing identical results across the original and the optimized model versions. The differences impacted the top-5 and top-10 predictions, while top-1 predictions remained unchanged. Although top-1 accuracy is often emphasized, variations in lower-ranked predictions can also reveal important issues. Our results show that focusing exclusively on top-1 accuracy may give developers a false sense that an optimization is successful, while overlooking substantial changes in top-K predictions (up to 2% in top-5 and 13% in top-10). Using OODTE, we detected these inconsistencies and reported them to the ONNX Optimizer developers.
@@
Table 3. Output label accuracy differences across classification (CL) and object detection (OD) models, showing type, opset version (Ver.), and accuracy gaps for top-1 (T-1), top-5 (T-5), and top-10 (T-10) predictions between original and ONNX-optimized models, along with the related issue (Is.): Format Version (FV) or Malformed Graph (MG).
# Model Type Ver. T-1 T-5 T-10 Is.
1 AlexNet (alexnet, ) CL 7-9 0% 0% 0.02% FV
2 DenseNet121 (densenet, ) CL 7-9 0% 0% 0.02% FV
3 GoogleNet (googlenet, ) CL 7-9 0% 0% 0.04% FV
4 Inception-1 (inception, ) CL 7-9 0% 0% 0.25% FV
5 Inception-2 (inception, ) CL 7-9 0% 0.02% 0.05% FV
6 MobileNet v2 (mobilenet, ; mobilenetv2, ) CL 7-9 0% 0% 0.02% FV
7 ResNet18 (resnet, ) CL 7 0% 0.02% 0.02% FV
8 ResNet34-v2 (resnetv2, ) CL 7 0% 0.02% 0.02% FV
9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
10 ShuffleNet-v1 (shufflenet, ) CL 7-9 0% 0.7% 1% FV
11 SqueezeNet 1.0 (squeezenet, ) CL 7-9 0% 0% 0.02% FV
12 EfficientNet-Lite4 (efficientnet, ) CL 11 0% 0.04% 1% MG
13 SSD (ssd, ) OD 12 0% 0% 0.04% MG
@@ 9 ResNet50-caffe2 (resnet, ) CL 7-9 0% 2% 13% FV
Implicit ONNX format version upgrade::
@@
The ONNX Optimizer silently upgrades models from ONNX format version v3 to v4, without any warning or indication. This behavior impacts models using opset versions 7, 8, and 9— that are older opsets but still supported in the current ONNX Runtime (v1.21.1). Despite being outdated, these opset versions are still prevalent, representing approximately 4141% of the models in the ONNX Model Hub (presented in Table 3, with the issue labeled as “FV”, indicating a problem with the format version.). While not all models within this opset range are affected, we observed a significant number that exhibited differences between their original and optimized versions. We have reported this issue to the ONNX Optimizer developers, who have acknowledged it and are investigating further.
Problematic fusion pass::
The pass fuse_bn_into_conv
fuses Batch Normalization and Convolution operators within the model. In particular, the pass attempts to fuse BatchNorma-
lization
@@
nodes (which do not update their mean or variance during inference) with their preceding Conv nodes by adjusting the weights and biases of the Conv node according to the properties of the BatchNormalization node. We observed that this pass was problematic in a number of cases, resulting in both a crashing bug and an issue where redundant nodes are introduced into the model. In the latter scenario, instead of properly merging the nodes resulting in a Conv node with adjusted weights and biases, the pass generates multiple nodes that represent the Batch Normalization formula. This behavior is illustrated on the right side of Figure 2. This formula representation is problematic as several models produced different output labels after optimization using this pass compared to their original versions (presented in Table 3, with the issue labeled as “MG”, indicating a malformed graph). Although another user previously reported a similar issue, no details on the reproduction steps or affected models were provided. We submitted a reproducible issue to the ONNX Optimizer developers, including all relevant information and behavior details.
@@
Table 3 shows label differences in top-K outputs for classification and object detection models, highlighting the causes.
@@
Table 4. Object detection metric differences between the original and optimized versions, using the default ONNX Optimizer setting. For each model, we present the average F1 score (F1), precision (Prec.), and recall (Rec.), as well as mAP and IoU.
Model Op. F1 Prec. Rec. mAP IoU
SSD (ssd, ) 12 99.99% 99.99% 99.99% 99.99% 100%
Tiny
YOLOv3 (Redmon2018YOLOv3AI, )
11 99.18% 99.18% 99.18% 99.99% 89.35%
YOLOv2 (yolov2, ) 9 99.95% 99.95% 99.95% 99.94% 88.49%
YOLOv3 (Redmon2018YOLOv3AI, ) 10 99.45% 99.45% 99.45% 99.99% 84.81%
YOLOv3 (Redmon2018YOLOv3AI, ) 12 99.45% 99.45% 99.45% 99.99% 84.81%
Figure 2. Incorrect model graph transformation caused by the fuse_bn_to_conv pass in EfficientNet-Lite-4. The nodes affected are highlighted. It should produce a Conv node with updated weights and biases, but the pass generates multiple nodes that represent the BatchNormalization formula.
Bounding Box Differences in Object Detection Models We observed some differences in the bounding box outputs between the original and optimized versions of object detection models. The results are presented in detail in Table 4 with the following main findings: 1. Both YOLOv3 models (opset $10$ and $12$) achieve high average IoU, F1 score, precision, and recall, all above 99.4599.45%. However, the optimized model occasionally misplaces bounding boxes—resulting in IoU drops of up to 15.1915.19%, these differences do not lead to misclassifications, as none were observed in the top ranked label predictions. A similar trend is also seen in YOLOv2 model where optimization causes a drop in IoU value but not in detection performance (F1, precision and recall).
6.3. Optimizer Warnings
@@
We observed two issues related to the addition of redundant elements (unused initializer entries), introduced in the optimized models, associated with the eliminate_shape_gather and eliminate_nop_reshape passes. When executed with ONNX Runtime, these entries were ignored, triggering a warning in the format: Removing initializer ‘INITIALIZER_ID’. It is not used by any node and should be removed from the model. Upon manual inspection, we found no significant differences in model size. However, this issue warrants further investigation for potential side effects towards model performance. We have reported it to the optimizer developers for additional review.
7. Tool Usability
@@
OODTE  can test whether the ONNX Optimizer generates optimized models without compromising output label correctness. Currently, OODTE supports differential testing for classification, object detection, and text generation models. However, OODTE can also be easily extended to support new model types, as long as (1) the models can be compiled into the ONNX format for optimization and (2) a comparator is provided to compare the results between the base and optimized models for the new model type. In addition, OODTE can fetch models and datasets automatically, making it adaptable for use in a Continuous Integration pipeline. This allows OODTE to test the ONNX Optimizer whenever new functionality (e.g., new passes) is added. By default, OODTE loads models from the official ONNX Model Hub (onnxmodelhub, ), which was used in our experiments. However, by configuring OODTE to use a custom ONNX Hub that follows the ONNX API specification for model repositories (onnxhubapi, ), it can fetch and use models without further modification. The same flexibility applies to datasets. OODTE can fetch datasets using packages like the datasets (datasets, ) package or from local storage.
@@
8. Threats To Validity
We list the threats to validity in our evaluation below:
Model Selection: Our evaluation covered 130 models from the official ONNX Model Hub, spanning classification, object detection, as well as text comprehension and generation. Models were chosen for their maturity, accessibility, and broad adoption. aimed to create a representative and diverse experimental set, although the selection is not exhaustive. We also excluded models with opset versions below 7 due to incompatibility with current ONNX and ONNX Runtime versions, limiting our results to newer models. To keep the study tractable, diffusion and reinforcement learning models were omitted; we plan to include these in future work.
Methodology Generalizability: We focus ONNX Optimizer because it is a core component of the widely used ONNX ecosystem (19.5K stars) (onnx, ) and supports tasks beyond model optimization, such as graph simplification (ONNX Simplifier, 4.1K stars (onnxsimplifier, )). Its popularity and broad applicability make it a strong candidate for study and a representative of the state-of-the-art technologies. While testing all DL compilers was infeasible, we demonstrated the value of our methodology and noted it can extend to other optimizers that offer per-pass APIs and runtime support for differential testing. For example, OODTE could support OpenVINO (openvino, ) and Apache TVM (tvm, ) by adapting their graph optimization passes, considering both compilers also provide runtime APIs, similarly to ONNX.
Optimization Pass Ordering:
The order of optimization passes can significantly affect model performance and behavior (OptimizationOrder, ), a well-known challenge in compiler research. While we do not study pass ordering in depth, OODTE adopts two evaluation modes: (1) the default ONNX Optimizer configuration, including common passes such as fuse and eliminate, and (2) a configuration applying all 47 available passes atomically. Although this does not capture issues from specific pass combinations or orderings, it effectively uncovers correctness problems in practice. Nevertheless, the omission of a systematic study on optimization ordering may limit the completeness of our evaluation.
9. Discussion